# Beyond the Hodge Theorem: curl and asymmetric pseudodifferential projections

### v3

@@ B.2 The spectrum
For a proof of the above theorem we refer the reader to [41, Theorem 5.5]. Somewhat more involuted proofs may also be found in [52, Lemma 4.1] and [38, Proposition 3.9].
Remark B.3.
When $a=1$ Theorem B.2 reduces to the usual spectrum of $\,-\Delta\,$ on the round 3-sphere: eigenvalues $n(n+2)$, $n=0,1,2,\dots$, with multiplicity $(n+1)^{2}\,$.
Appendix C The spectrum of curl on a Berger sphere
In this appendix we will write down explicitly the eigenvalues and the $\eta$-invariant for the operator curl on a Berger sphere. This will be done using classical (as opposed to pseudodifferential) techniques and will serve as an illustration of the spectral asymmetry of curl.
Throughout this appendix we use the notation from Appendix B.1.
C.1 The spectrum
The spectrum of curl (B.7) was first computed by Gibbons [29, Section 5]. The positive spectrum of curl (B.7) also features in [41, Proposition 5.6]. We give below, for future reference and in the form of a theorem, explicit formulae for the full spectrum of curl, in a slightly reformulated form. The proof is omitted.
Theorem C.1.
@@ C.1 The spectrum
    $\frac{n}{a}\,,\qquad n=2,3,\dots,$ (C.1)
@@
    $\frac{n+2(a^{2}-1)}{a}\,,\qquad n=2,3,\dots.$ (C.3)
@@
    $a+\sqrt{a^{2}+n(n+2)+\left(a^{-2}-1\right)\left(n-2l\right)^{2}}\ ,\quad n=2,3,\dots,\quad 1\leq l\leq\left\lfloor\frac{n}{2}\right\rfloor.$ (C.6)
@@
    $a-\sqrt{a^{2}+n(n+2)+\left(a^{-2}-1\right)\left(n-2l\right)^{2}}\ ,\quad n=2,3,\dots,\quad 1\leq l\leq\left\lfloor\frac{n}{2}\right\rfloor,$ (C.10)
    with multiplicities as in (C.7)–(C.9).
Remark C.2.
@@
    For $a=1$ the Berger sphere reduces to the standard 3-sphere. It is not hard to check that for $a=1$ Theorem C.1 gives us eigenvalues
    ±$\pm n,\qquad n=2,3,\dots,$ (C.11)
@@
    $n^{2}-1.$ (C.12)
    In particular, the spectrum is symmetric about 00.
@@
    Formulae (C.11) and (C.12) agree with [7, Theorem 5.2].
@@
    Note that in (C.6) the lowest value the integer $l$ can take is $1$. The case $l=0$ is special: in this case the square root becomes superfluous and the expression (C.6) turns into (C.3). We have highlighted the special status of the case $l=0$ by separating out the series of eigenvalues (C.3).
@@
    Observe that the expression (B.8) appears under the square root in formulae (C.6) and (C.10).
C.2 The eta invariant
In this appendix we will compute the eta invariant for the operator $\operatorname{curl}$ on a Berger 3-sphere directly, using the explicit formulae for the eigenvalues established above in Theorem C.1. Recall that the eta-function is defined, in general, in accordance with (1.2).
Theorem C.3.
Consider the operator $\operatorname{curl}$ on a Berger 3-sphere, defined as in Appendix B.1. We have
$\eta_{\operatorname{curl}}(0)=\frac{2}{3}(a^{2}-1)^{2}.$ (C.13)
Remark C.4.
@@ C.2 The eta invariant
    The eta invariant for the Dirac operator on a Berger 3-sphere was computed by Hitchin [32, p. 34] and reads
    $\eta_{\mathrm{Dirac}}(0)=-\frac{1}{6}(a^{2}-1)^{2}\,.$ (C.14)
    Note that the parameter $\lambda$ in [32] corresponds to our parameter $a$, and that Hitchin adopts the same (or rather, equivalent) conventions as we do when defining the Dirac operator. The latter can be seen if one compares the spectra of the Dirac operator in [32] and [27]: Hitchin’s [32, Proposition 3.2] with $p=q=1$ and minus sign in front of the square root agrees with [27, formulae (6.5) and (6.8)].
    The result (C.14) was obtained under the assumption $a<4$. We do not need to impose any restriction on $a$ because, unlike for Dirac operator, where the dimension of the space of harmonic spinors is highly dependent on the Riemannian metric, zero is never an eigenvalue of curl (see Theorem 2.1).
@@
    $\eta_{\operatorname{curl}}(0)=-4\eta_{\mathrm{Dirac}}(0)+k$ (C.15)
    for some $k\in\mathbb{Z}$. This follows from abstract algebraic topology arguments involving Hirzebruch polynomials, see [2, Section 1] and [3, Theorems 4.2 and 4.14]. Theorem C.3 establishes that, in the case of a Berger 3-sphere, we have (C.15) with $k=0$ for $a<4$. The significance of the value $a=4$ is that for this particular value of the parameter $a$ the Dirac operator has a zero mode (harmonic spinor).
@@
    Formula (C.13) appears in [23, Problem 7].
In preparation for the proof of Theorem C.3, let us make the following observation.
Lemma C.5.
@@
$\eta_{\operatorname{curl}}(s)=\theta(s)+(2a)^{-s}+4a^{s}\zeta(s-1)\,,$ (C.16)
@@
$\theta(s):=\sum_{j=1}^{\infty}\left[\left(\sqrt{a^{2}+\mu_{j}}\,+\,a\right)^{-s}-\left(\sqrt{a^{2}+\mu_{j}}\,-\,a\right)^{-s}\right],$ (C.17)
the $\mu_{j}$ are the eigenvalues (2.17) of $\,-\Delta\,$ enumerated in increasing order with account of multiplicities, and $\,\zeta(s):=\sum_{n=1}^{\infty}n^{-s}\,$ is the Riemann zeta function.
@@
Formulae (C.16) and (C.17) are a straightforward consequence of Theorems B.2 and C.1. ∎
Definition C.6.
@@
to be the zeta function [^5] The function $s\mapsto\zeta_{-\Delta}(s)$ is often called the Minakshisundaram–Pleijel zeta function, see [44, equation (6)]. associated with (minus) the Laplace–Beltrami operator $-\Delta$. Similarly, we define
@@
It is known [44, p. 243] that the function $\zeta_{\sqrt{-\Delta}}(s)$ is meromorphic in $\mathbb{C}$ with simple poles at $s=3,1,-1,\dots$.
Lemma C.7.
@@
$\zeta_{\sqrt{-\Delta}}(s)=\frac{1}{2\pi^{2}}\left[\frac{\operatorname{Vol}(M)}{s-3}+\frac{\int_{M}\operatorname{Sc}(x)\,\rho(x)\,\mathrm{d}x}{12(s-1)}+\ldots\right].$ (C.18)
@@
Let $N_{\sqrt{-\Delta}}(\lambda)$ be the counting function for the positive eigenvalues of $\sqrt{-\Delta}$, see, e.g., [14, Appendix B]. The functions $\zeta_{\sqrt{-\Delta}}$ and $N_{\sqrt{-\Delta}}$ are related as
$\zeta_{\sqrt{-\Delta}}(s)=\int_{0}^{+\infty}\lambda^{-s}\,N^{\prime}_{\sqrt{-\Delta}}(\lambda)\,\mathrm{d}\lambda\,,$ (C.19)
where the prime stands for differentiation with respect to $\lambda$. By combining [14, equation (B.1) and Theorem B.2] with formula (C.19), we obtain (C.18). ∎
@@
$\zeta_{\sqrt{-\Delta}}(s)=a\left[\frac{1}{s-3}+\frac{4-a^{2}}{6(s-1)}+\ldots\right].$ (C.20)
We are now in a position to prove Theorem C.3.
Proof of Theorem C.3.
Since $\,\zeta(-1)=-\frac{1}{12}\,$, formula (C.16) implies that the proof of (C.13) reduces to the proof of
$\theta(0)=\frac{2}{3}\,a^{2}(a^{2}-2)\,,$ (C.21)
where $\theta$ is the function (C.17). Formulae (C.21) and (C.17) reduce the analysis of the operator $\,\operatorname{curl}\,$ to the analysis of the Laplace–Beltrami operator.
To establish (C.21) we use a trick due to Hitchin [32, p. 34]. We expand the terms in the RHS of (C.17) in inverse powers of $\mu_{j}$ as
$\Bigl{(}\sqrt{a^{2}+\mu_{j}}\,\pm\,a\Bigr{)}^{-s}=\mu_{j}^{-s/2}\Bigl{(}\sqrt{1+a^{2}\mu_{j}^{-1}}\,\pm\,a\mu_{j}^{-1/2}\Bigr{)}^{-s}\\ =\mu_{j}^{-s/2}\Bigl{(}1\pm a\mu_{j}^{-1/2}+\frac{1}{2}a^{2}\mu_{j}^{-1}+O\left(\mu_{j}^{-2}\right)\Bigr{)}^{-s}\\ =\mu_{j}^{-s/2}\Bigl{(}1-s\Bigl{[}\pm a\mu_{j}^{-1/2}+\frac{1}{2}a^{2}\mu_{j}^{-1}\Bigr{]}+\frac{s(s+1)}{2}\Bigl{[}a^{2}\mu_{j}^{-1}\pm a^{3}\mu_{j}^{-3/2}\Bigr{]}\mp\frac{s(s+1)(s+2)}{6}a^{3}\mu_{j}^{-3/2}+O\left(\mu_{j}^{-2}\right)\Bigr{)}\\ =\mu_{j}^{-s/2}\Bigl{(}1\mp sa\mu_{j}^{-1/2}+\frac{s^{2}}{2}a^{2}\mu_{j}^{-1}\pm\frac{s(1-s^{2})}{6}a^{3}\mu_{j}^{-3/2}+O\left(\mu_{j}^{-2}\right)\Bigr{)}.$ (C.22)
Substituting (C.22) into (C.17) we get
[…]
@@
$h(0)=0\,.$ (C.24)
Formulae (C.23), (C.20) and (C.24) imply
[…]
which proves (C.21). ∎
Appendix D Maxwell’s equations and electromagnetic chirality
The focus of our paper is the study of asymmetry between positive and negative eigenvalues of the operator $\operatorname{curl}$. It is natural to ask the question: does the sign of an eigenvalue of $\operatorname{curl}$ have a physical meaning? We show in this appendix that it does. The key here is the notion of electromagnetic chirality, see Definition D.3.
@@ Appendix D Maxwell’s equations and electromagnetic chirality
$\begin{pmatrix}0&\operatorname{curl}\\ -\operatorname{curl}&0\end{pmatrix}\begin{pmatrix}E\\ B\end{pmatrix}=\frac{\partial}{\partial t}\begin{pmatrix}E\\ B\end{pmatrix},$ (D.1)
$\delta E=\delta B=0,$ (D.2)
@@
We list below two basic symmetries of Maxwell’s equations (D.1), (D.2). All the results in this appendix are given without proof because the corresponding proofs are elementary.
Definition D.1.
@@
$\operatorname{dual}:\begin{pmatrix}E\\ B\end{pmatrix}\mapsto\begin{pmatrix}-B\\ E\end{pmatrix}.$ (D.3)
@@
$F_{\alpha\beta}=\begin{pmatrix}0&E_{1}&E_{2}&E_{3}\\ -E_{1}&0&-B_{3}&B_{2}\\ -E_{2}&B_{3}&0&-B_{1}\\ -E_{3}&-B_{2}&B_{1}&0\end{pmatrix}.$ (D.4)
@@
Lemma D.2.
Maxwell’s equations (D.1), (D.2) are invariant under the action of the duality transform (D.3), i.e. the duality transform maps solutions to solutions.
Definition D.3 ([51, 40]).
@@
$\operatorname{chir}(E,B):=\int_{M}(E\wedge\mathrm{d}E+B\wedge\mathrm{d}B).$ (D.5)
Lemma D.4.
Chirality (D.5) is a conserved quantity for Maxwell’s equations (D.1), (D.2), i.e. if $(E,B)$ is a solution then $\operatorname{chir}(E,B)$ does not depend on time $t$.
@@
Definition D.5.
A harmonic solution $(E,B)$ of Maxwell’s equations (D.1), (D.2) is said to be polarised if $\,(\dot{E},\dot{B})=\pm\lambda\operatorname{dual}(E,B)\,$. Here the dot stands for derivative in $t$ and $\lambda$ is the angular frequency.
Recall that $(\lambda_{j},u_{j})$, $j=\pm 1,\pm 2,\ldots$, is the eigensystem for $\operatorname{curl}$. It is easy to see that
$\begin{pmatrix}E\\ B\end{pmatrix}=\begin{pmatrix}u_{j}\cos(\lambda_{j}t)\\ -u_{j}\sin(\lambda_{j}t)\end{pmatrix}$ (D.6)
@@
$\begin{pmatrix}E\\ B\end{pmatrix}=\begin{pmatrix}u_{j}\sin(\lambda_{j}t)\\ u_{j}\cos(\lambda_{j}t)\end{pmatrix}$ (D.7)
are harmonic polarised solutions of Maxwell’s equations (D.1), (D.2). The two solutions (D.6) and (D.7) are related via the duality transform (D.3) and describe waves with opposite polarisations. We see that the correspondence between eigenvalues and eigenforms of $\operatorname{curl}$ on the one hand and harmonic polarised solutions of Maxwell’s equations on the other is one-to-two.
@@
$\left.E\right|_{t=0}=E_{0},\qquad\left.B\right|_{t=0}=B_{0}$ (D.8)
for Maxwell’s equations (D.1), (D.2). It is easy to see that the solution can be written as a series
$\begin{pmatrix}E\\ B\end{pmatrix}=\sum_{j\in\mathbb{Z}\setminus\{0\}}\left[\begin{pmatrix}u_{j}\cos(\lambda_{j}t)\\ -u_{j}\sin(\lambda_{j}t)\end{pmatrix}\langle E_{0}\,,u_{j}\rangle+\begin{pmatrix}u_{j}\sin(\lambda_{j}t)\\ u_{j}\cos(\lambda_{j}t)\end{pmatrix}\langle B_{0}\,,u_{j}\rangle\right]+\begin{pmatrix}P_{\mathcal{H}^{1}}E_{0}\\ P_{\mathcal{H}^{1}}B_{0}\end{pmatrix},$ (D.9)
[…]
@@
Proposition D.6.
For both solutions (D.6) and (D.7) we have
$\operatorname{chir}(E,B)=\lambda_{j}\,.$ (D.10)
Formula (D.10) tells us that the sign of an eigenvalue of $\operatorname{curl}$ has the physical meaning of the sign of chirality of the corresponding harmonic polarised solution of Maxwell’s equations. And here it does not matter which polarisation one chooses when writing down the harmonic solution of Maxwell’s equations.
Appendix E Taylor expansions for the operator of parallel transport

### v4

@@ C.2 The spectrum
For a proof of the above theorem we refer the reader to [50, Theorem 5.5]. Somewhat more involuted proofs may also be found in [62, Lemma 4.1] and [45, Proposition 3.9].
Remark C.3.
When $a=1$ Theorem C.2 reduces to the usual spectrum of $\Delta\,$ on the round 3-sphere: eigenvalues $n(n+2)$, $n=0,1,2,\dots$, with multiplicity $(n+1)^{2}\,$.
Appendix D The spectrum of curl on a Berger sphere
From a perturbative perspective, and in analogy with what happens for the massless Dirac operator, one expects the spectrum of curl to be asymmetric about zero for an open dense subset of the space of metrics on a given Riemannian 3-manifold — put differently, spectral asymmetry of curl is a generic condition. However, we were unable to identify a precise reference for this fact in the literature.
In this appendix we will write down explicitly the eigenvalues and the $\eta$-invariant for the operator curl on a Berger sphere, using classical (as opposed to pseudodifferential) techniques. On the one hand, this will serve as an illustration of the spectral asymmetry of curl. On the other hand, it shall serve as a useful reference for further explicit calculations in future work.
Throughout this appendix we use the notation from Appendix C.1.
D.1 The spectrum
The spectrum of curl (C.7) was first computed by Gibbons [34, Section 5]. The positive spectrum of curl (C.7) also features in [50, Proposition 5.6]. We give below, for future reference and in the form of a theorem, explicit formulae for the full spectrum of curl, in a slightly reformulated form. The proof is omitted.
Theorem D.1.
@@ D.1 The spectrum
    $\frac{n}{a}\,,\qquad n=2,3,\dots,$ (D.1)
@@
    $\frac{n+2(a^{2}-1)}{a}\,,\qquad n=2,3,\dots.$ (D.3)
@@
    $a+\sqrt{a^{2}+n(n+2)+\left(a^{-2}-1\right)\left(n-2l\right)^{2}}\ ,\quad n=2,3,\dots,\quad 1\leq l\leq\left\lfloor\frac{n}{2}\right\rfloor.$ (D.6)
@@
    $a-\sqrt{a^{2}+n(n+2)+\left(a^{-2}-1\right)\left(n-2l\right)^{2}}\ ,\quad n=2,3,\dots,\quad 1\leq l\leq\left\lfloor\frac{n}{2}\right\rfloor,$ (D.10)
    with multiplicities as in (D.7)–(D.9).
Remark D.2.
@@
    For $a=1$ the Berger sphere reduces to the standard 3-sphere. It is not hard to check that for $a=1$ Theorem D.1 gives us eigenvalues
    $\pm n,\qquad n=2,3,\dots,$ (D.11)
@@
    $n^{2}-1.$ (D.12)
    In particular, the spectrum is symmetric about 0.
@@
    Formulae (D.11) and (D.12) agree with [7, Theorem 5.2].
@@
    Note that in (D.6) the lowest value the integer $l$ can take is $1$. The case $l=0$ is special: in this case the square root becomes superfluous and the expression (D.6) turns into (D.3). We have highlighted the special status of the case $l=0$ by separating out the series of eigenvalues (D.3).
@@
    Observe that the expression (C.8) appears under the square root in formulae (D.6) and (D.10).
D.2 The eta invariant
In this subsection we will compute the eta invariant for the operator $\operatorname{curl}$ on a Berger 3-sphere directly, using the explicit formulae for the eigenvalues established above in Theorem D.1. Recall that the eta-function is defined, in general, in accordance with (1.2).
Theorem D.3.
Consider the operator $\operatorname{curl}$ on a Berger 3-sphere, defined as in Appendix C.1. We have
$\eta_{\operatorname{curl}}(0)=\frac{2}{3}(a^{2}-1)^{2}.$ (D.13)
Remark D.4.
@@ D.2 The eta invariant
    The eta invariant for the Dirac operator on a Berger 3-sphere was computed by Hitchin [37, p. 34] and reads
    $\eta_{\mathrm{Dirac}}(0)=-\frac{1}{6}(a^{2}-1)^{2}\,.$ (D.14)
    Note that the parameter $\lambda$ in [37] corresponds to our parameter $a$, and that Hitchin adopts the same (or rather, equivalent) conventions as we do when defining the Dirac operator. The latter can be seen if one compares the spectra of the Dirac operator in [37] and [31]: Hitchin’s [37, Proposition 3.2] with $p=q=1$ and minus sign in front of the square root agrees with [31, formulae (6.5) and (6.8)].
    The result (D.14) was obtained under the assumption $a<4$. We do not need to impose any restriction on $a$ because, unlike for Dirac operator, where the dimension of the space of harmonic spinors is highly dependent on the Riemannian metric, zero is never an eigenvalue of curl (see Theorem 2.1).
@@
    $\eta_{\operatorname{curl}}(0)=-4\eta_{\mathrm{Dirac}}(0)+k$ (D.15)
    for some $k\in\mathbb{Z}$. This follows from abstract algebraic topology arguments involving Hirzebruch polynomials, see [2, Section 1] and [3, Theorems 4.2 and 4.14]. Theorem D.3 establishes that, in the case of a Berger 3-sphere, we have (D.15) with $k=0$ for $a<4$. The significance of the value $a=4$ is that for this particular value of the parameter $a$ the Dirac operator has a zero mode (harmonic spinor).
@@
    Formula (D.13) appears in [25, Problem 7].
In preparation for the proof of Theorem D.3, let us make the following observation.
Lemma D.5.
@@
$\eta_{\operatorname{curl}}(s)=\theta(s)+(2a)^{-s}+4a^{s}\zeta(s-1)\,$, (D.16)
@@
$\theta(s):=\sum_{j=1}^{\infty}\left[\left(\sqrt{a^{2}+\mu_{j}}\,+\,a\right)^{-s}-\left(\sqrt{a^{2}+\mu_{j}}\,-\,a\right)^{-s}\right],$ (D.17)
the $\mu_{j}$ are the eigenvalues (B.12) of $\Delta\,$ enumerated in increasing order with account of multiplicities, and $\zeta(s):=\sum_{n=1}^{\infty}n^{-s}\,$ is the Riemann zeta function.
@@
Formulae (D.16) and (D.17) are a straightforward consequence of Theorems C.2 and D.1. ∎
Definition D.6.
@@
to be the zeta function [^5] The function $s\mapsto\zeta_{-\Delta}(s)$ is often called the Minakshisundaram–Pleijel zeta function, see [54, equation (6)]. associated with (minus) the Laplace–Beltrami operator $\Delta$. Similarly, we define
@@
It is known [54, p. 243] that the function $\zeta_{\sqrt{-\Delta}}(s)$ is meromorphic in $\mathbb{C}$ with simple poles at $s=3,1,-1,\dots$.
Lemma D.7.
@@
$\zeta_{\sqrt{-\Delta}}(s)=\frac{1}{2\pi^{2}}\left[\frac{\operatorname{Vol}(M)}{s-3}+\frac{\int_{M}\operatorname{Sc}(x)\,\rho(x)\,\mathrm{d}x}{12(s-1)}+\ldots\right].$ (D.18)
@@
Let $N_{\sqrt{-\Delta}}(\lambda)$ be the counting function for the positive eigenvalues of $\sqrt{-\Delta}$, see, e.g., [15, Appendix B]. The functions $\zeta_{\sqrt{-\Delta}}$ and $N_{\sqrt{-\Delta}}$ are related as
$\zeta_{\sqrt{-\Delta}}(s)=\int_{0}^{+\infty}\lambda^{-s}\,N^{\prime}_{\sqrt{-\Delta}}(\lambda)\,\mathrm{d}\lambda\,$, (D.19)
where the prime stands for differentiation with respect to $\lambda$. By combining [15, equation (B.1) and Theorem B.2] with formula (D.19), we obtain (D.18). ∎
@@
$\zeta_{\sqrt{-\Delta}}(s)=a\left[\frac{1}{s-3}+\frac{4-a^{2}}{6(s-1)}+\ldots\right].$ (D.20)
We are now in a position to prove Theorem D.3.
Proof of Theorem D.3.
Since $\zeta(-1)=-\frac{1}{12}\,$, formula (D.16) implies that the proof of (D.13) reduces to the proof of
$\theta(0)=\frac{2}{3}\,a^{2}(a^{2}-2)\,$, (D.21)
where $\theta$ is the function (D.17). Formulae (D.21) and (D.17) reduce the analysis of the operator $\operatorname{curl}\,$ to the analysis of the Laplace–Beltrami operator.
To establish (D.21) we use a trick due to Hitchin [37, p. 34]. We expand the terms in the RHS of (D.17) in inverse powers of $\mu_{j}$ as
$\Bigl(\sqrt{a^{2}+\mu_{j}}\,\pm\,a\Bigr)^{-s}=\mu_{j}^{-s/2}\Bigl(\sqrt{1+a^{2}\mu_{j}^{-1}}\,\pm\,a\mu_{j}^{-1/2}\Bigr)^{-s}\\ =\mu_{j}^{-s/2}\Bigl(1\pm a\mu_{j}^{-1/2}+\frac{1}{2}a^{2}\mu_{j}^{-1}+O\left(\mu_{j}^{-2}\right)\Bigr)^{-s}\\ =\mu_{j}^{-s/2}\Bigl(1-s\Bigl[\pm a\mu_{j}^{-1/2}+\frac{1}{2}a^{2}\mu_{j}^{-1}\Bigr]+\frac{s(s+1)}{2}\Bigl[a^{2}\mu_{j}^{-1}\pm a^{3}\mu_{j}^{-3/2}\Bigr]\mp\frac{s(s+1)(s+2)}{6}a^{3}\mu_{j}^{-3/2}+O\left(\mu_{j}^{-2}\right)\Bigr)\\ =\mu_{j}^{-s/2}\Bigl(1\mp sa\mu_{j}^{-1/2}+\frac{s^{2}}{2}a^{2}\mu_{j}^{-1}\pm\frac{s(1-s^{2})}{6}a^{3}\mu_{j}^{-3/2}+O\left(\mu_{j}^{-2}\right)\Bigr).$ (D.22)
Substituting (D.22) into (D.17) we get
$\theta(s)=-\,2sa\,\zeta_{\sqrt{-\Delta}}(s+1)\,+\,\frac{s(1-s^{2})}{3}\,a^{3}\,\zeta_{\sqrt{-\Delta}}(s+3)\,+\,h(s)\,$, (D.23)
@@
h(0)=0.h(0)=0\,. (D.24)
Formulae (D.23), (D.20) and (D.24) imply
$\lim_{s\to 0}\theta(s)=-\,2a^{2}\,\frac{4-a^{2}}{6}\,+\,\frac{1}{3}\,a^{4}=\frac{2}{3}\,a^{2}(a^{2}-2)\,$,
which proves (D.21). ∎
Appendix E Maxwell’s equations: electromagnetic chirality and polarisation
The focus of our paper is the study of asymmetry between positive and negative eigenvalues of the operator $\operatorname{curl}$. It is natural to ask the question: does the sign of an eigenvalue of $\operatorname{curl}$ have a physical meaning? We show in this appendix that it does by means of a rather elementary argument. The key here are the notions of electromagnetic chirality and polarisation, see Definitions E.3 and E.5. Let us emphasise that the physical interpretation we provide in this appendix is not new: it has appeared in the literature in various related, if more abstract and ‘higher-tech’, formulations — see, e.g., [30, Sec. 2], [33, Sec. 1.1.2] and references therein.
@@ Appendix E Maxwell’s equations: electromagnetic chirality and polarisation
$\begin{pmatrix}0&\operatorname{curl}\\ -\operatorname{curl}&0\end{pmatrix}\begin{pmatrix}E\\ B\end{pmatrix}=\frac{\partial}{\partial t}\begin{pmatrix}E\\ B\end{pmatrix}$, (E.1)
$\delta E=\delta B=0,$ (E.2)
@@
We list below two basic symmetries of Maxwell’s equations (E.1), (E.2). All the results in this appendix are given without proof because the corresponding proofs are elementary.
Definition E.1.
@@
$\operatorname{dual}:\begin{pmatrix}E\\ B\end{pmatrix}\mapsto\begin{pmatrix}-B\\ E\end{pmatrix}$. (E.3)
@@
$F_{\alpha\beta}=\begin{pmatrix}0&E_{1}&E_{2}&E_{3}\\ -E_{1}&0&-B_{3}&B_{2}\\ -E_{2}&B_{3}&0&-B_{1}\\ -E_{3}&-B_{2}&B_{1}&0\end{pmatrix}$.
@@
Lemma E.2.
Maxwell’s equations (E.1), (E.2) are invariant under the action of the duality transform (E.3), i.e. the duality transform maps solutions to solutions.
Definition E.3 ([61, 49]).
@@
$\operatorname{chir}(E,B):=\int_{M}(E\wedge\mathrm{d}E+B\wedge\mathrm{d}B).$ (E.4)
Lemma E.4.
Chirality (E.4) is a conserved quantity for Maxwell’s equations (E.1), (E.2), i.e. if $(E,B)$ is a solution then $\operatorname{chir}(E,B)$ does not depend on time $t$.
@@
Definition E.5.
A harmonic solution $(E,B)$ of Maxwell’s equations (E.1), (E.2) is said to be polarised if $(\dot{E},\dot{B})=\pm\lambda\operatorname{dual}(E,B)\,$. Here the dot stands for derivative in $t$ and $\lambda$ is the angular frequency.
Recall that $(\lambda_{j},u_{j})$, $j=\pm 1,\pm 2,\ldots$, is the eigensystem for $\operatorname{curl}$. Consider a harmonic solution
$\begin{pmatrix}E\\ B\end{pmatrix}=\begin{pmatrix}u_{j}\cos(\lambda_{j}t+\varphi)\\ -u_{j}\sin(\lambda_{j}t+\varphi)\end{pmatrix}$ (E.5)
of Maxwell’s equations (E.1), (E.2), where $\varphi\in\mathbb{R}$ is an arbitrary phase. Of course, (E.5) is a linear combination of two basic harmonic solutions
$\begin{pmatrix}E\\ B\end{pmatrix}=\begin{pmatrix}u_{j}\cos(\lambda_{j}t)\\ -u_{j}\sin(\lambda_{j}t)\end{pmatrix}$ (E.6)
@@
$\begin{pmatrix}E\\ B\end{pmatrix}=\begin{pmatrix}u_{j}\sin(\lambda_{j}t)\\ u_{j}\cos(\lambda_{j}t)\end{pmatrix}$. (E.7)
@@
$\left.E\right|_{t=0}=E_{0},\qquad\left.B\right|_{t=0}=B_{0}$
for Maxwell’s equations (E.1), (E.2). Then, (E.6) and (E.7) can be used to write the solution as a series
$\begin{pmatrix}E\\ B\end{pmatrix}=\sum_{j\in\mathbb{Z}\setminus\{0\}}\left[\begin{pmatrix}u_{j}\cos(\lambda_{j}t)\\ -u_{j}\sin(\lambda_{j}t)\end{pmatrix}\langle E_{0}\,,u_{j}\rangle+\begin{pmatrix}u_{j}\sin(\lambda_{j}t)\\ u_{j}\cos(\lambda_{j}t)\end{pmatrix}\langle B_{0}\,,u_{j}\rangle\right]+\begin{pmatrix}P_{\mathcal{H}^{1}}E_{0}\\ P_{\mathcal{H}^{1}}B_{0}\end{pmatrix}$,
where $P_{\mathcal{H}^{1}}$ is the orthogonal projection onto the space of harmonic 1-forms and $\langle\,\cdot\$$\cdot\,\rangle$ is the inner product (2.5).
@@
Proposition E.6.
For (E.5) we have
$\operatorname{chir}(E,B)=\lambda_{j}\,,\qquad\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\dot{E},\dot{B})=-\lambda_{j}\operatorname{dual}(E,B).\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$ (E.8)
Formula (E.8) tells us that the sign of an eigenvalue of $\operatorname{curl}$ has a physical meaning and it can be interpreted in two different ways — in terms of chirality of the corresponding harmonic solution (E.5) of Maxwell’s equations or in terms of its polarisation. Note that the concepts of chirality and polarisation are different: polarisation involves time, whereas chirality does not.
Appendix F Taylor expansions for the operator of parallel transport